OpenStudy (anonymous):

show that y=tant satisfies the IVP y'=1/(1+t^2), y(0)=0. what is the largest open interval containing t=0 over which y=tant is a solution?

OpenStudy (anonymous):

i am confused because integral of 1/(1+t^2) is arctant...

OpenStudy (amistre64):

if y=tan(t) IS a solution then itll fit in just nicely

OpenStudy (anonymous):

ok, cuz it actually if a solution...but shouldn't the question have assigned different variables??

OpenStudy (anonymous):

*is a solution

OpenStudy (amistre64):

no, t is t is t in all the derivatives of y

OpenStudy (anonymous):

well how do i find the limit of t now? both y and y' have to considered right? so i can find the largest interval?

OpenStudy (amistre64):

well, tan(t) is that has a zero in it, and is contiuous would be: (-pi/2 , pi/2) if i see this right

OpenStudy (anonymous):

right

OpenStudy (amistre64):

you may be reading to much into the problem :)

OpenStudy (turingtest):

but y=tan^(-1)t...

OpenStudy (turingtest):

though the interval -pi/2,pi/2 still makes sense

OpenStudy (anonymous):

i think so too cuzi still don't understand... so limit of t in y' would be neg inf to pos inf but y is restricted so in the end, largest interval is -pi/2 to pi/2?

OpenStudy (turingtest):

arctan(t) is undefined at \[\pi/2+n\pi\]

OpenStudy (amistre64):

is y=tan(t) a solution?

OpenStudy (turingtest):

no, y'=sec^2t

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok, no

OpenStudy (amistre64):

or is it; y=arctan(t) a solution? are you reading it correctly is the quandry

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=%20y'%3D1%2F(1%2Bt%5E2)&t=crmtb01 in case you didn't know the derivative of arctan...

OpenStudy (anonymous):

i do know the derivative that's why i was getting confused. cuz the y in the prolem doesn't equal integral of y'. ok so no, not a solution...so limit IS neg inf to pos inf

OpenStudy (amistre64):

as is, we are stating: \[tan'(t)=\frac{1}{1+t^2}\] \[sec^2(t)=\frac{1}{1+t^2}\]

OpenStudy (turingtest):

no I said earlier it should be arctan tan is not a solution

OpenStudy (amistre64):

if not a solution; then there is no interval to worry about since no interval is a solution to a nonsolution :)

OpenStudy (anonymous):

right

OpenStudy (turingtest):

right!

OpenStudy (anonymous):

but arctan(0)=0

OpenStudy (amistre64):

\[\frac{dy}{dt}=\frac{1}{1+t^2}\] \[{dy}=\frac{1}{1+t^2}dt\] \[\int\ ({dy}=\frac{1}{1+t^2}dt)\] \[y=tan^{-1}(t)+C_1\]

OpenStudy (anonymous):

mmhmm

OpenStudy (anonymous):

C=0 so y=arctant

OpenStudy (amistre64):

when y=0 and t=0; we get: \[0=tan^{-1}(0)+C_1\] \[0=C_1\] therefore, given the IVP \[y=tan^{-1}(t)\]

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

this shows that y=tan(t) is NOT a solution to the IVP and is therefore no need to worry about any nonexisting intervals

OpenStudy (anonymous):

ok thank you so much